3500 2pF 210mH 10000

Answer the following with all steps, thanks

For the circuit shown, ig= 50 cos 20,000t mA

1. Find the inductive and capacitive reactances

2. Construct the phasor domain equivalent circuit

3. Find the individual branch impedances

4. Find the current I through the capacitive branch if ig = 50 cos 20,000t mA

5. Find the steady-state expression for voltage vo
Answer #1

Son 2uF 3 JOMH 1001 River, - 1) In 1q = 50 cos 20000t ma. phasor form, Ig = 5o Loo ma & w=20000 rad /sec. 1. Inductive Reacta3-) let Equivalent branch impedances be a,,22 BI sor, L in series > 50+ z = (50 + 9 200) i 2. = 50 + 320012 coor in series >